OpenStudy (anonymous):

three capacitors of capacitances 6meuF each are available.the minimum and the maximum capacitances ,which may be obtained are a 6,18 b 3,12 c 2,12 d 2,18

OpenStudy (abhisar):

Hello @fruits ! \(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\sf\color{blue}{~~~~Welcome~to~OpenStudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

OpenStudy (abhisar):

Minimum will be when we group the three capacitances into series combo. so the min will be 2 and max will be when we group them in parallel, so max will be 18

OpenStudy (abhisar):

In series combo \[\frac{ 1 }{ Cf }=\frac{ 1 }{ C1 }+\frac{ 1 }{ C2 }+ .......\]

OpenStudy (abhisar):

In parallel combo, \[C_{f}=C_{1}+C_{2}+......\]

OpenStudy (anonymous):

ok thank you

OpenStudy (abhisar):

\(\Huge\text{Anytime !}\) \(\huge\ddot\smile\)

OpenStudy (anonymous):

can i know in which standard u are in now??

OpenStudy (abhisar):

http://openstudy.com/users/abhisar Here u will find all the info

OpenStudy (abhisar):

\(\color{green}{\huge\ddot\smile}\)

OpenStudy (anonymous):

ohh so great:D

OpenStudy (abhisar):

u too \(\color{green}{\huge\ddot\smile}\)

OpenStudy (anonymous):

M still in 12th

OpenStudy (abhisar):

That's also not easy :D

OpenStudy (anonymous):

hmm u r right:d

OpenStudy (abhisar):

All the best for ur Future ! \(\color{green}{\huge\ddot\smile}\)

OpenStudy (anonymous):

thanx